OpenStudy (anonymous):

An object is dropped from a height of 1700 ft above the ground. The function h=-16t^2+1700 gives the object’s height h in feet during free fall at t seconds. a. When will the object be at 1000ft above the ground? b. When will the object be 940ft above the ground? c. What are a reasonable domain and range for the function h?

OpenStudy (cruffo):

both part a and b are asking you to find the time t for a given h. for a, solve the following equations for t: -16t^2 + 1700 = 1000

OpenStudy (anonymous):

t=(close to) 1.654 is what I got a.

OpenStudy (cruffo):

I got t = 6.614 -16t^2 +1700 = 1000 -16t^2 = -700 t^2 = 43.75 then take square root of both sides

OpenStudy (anonymous):

1000=-16t^2+1700 -700=-16t^2 _/-700=_/-16t^2 26.458/16=16t/16 t=1.654 is what I did

OpenStudy (cruffo):

You need to divide by 16 first, before you take the square root.

OpenStudy (anonymous):

Okay got a. and b. now I just need c.:)

OpenStudy (cruffo):

Domain = what values of time would be used here? t > 0, since negative time not reasonable. And you probably need to find when the ball hits the ground, since after that time, the ball would be underground.

OpenStudy (anonymous):

1700=16t^2 _/-106.25=t^2 t=10.3078 Is what I got for how long it takes to hit the ground.

OpenStudy (cruffo):

Cool. So the a reasonable domain is from 0 to 10.3078.

OpenStudy (cruffo):

Range: what values of height seem reasonable? h > 0, because negative height would mean that the ball was underground. And ... since the ball was "dropped" from a height of 1700 (vs being "thrown up" from a height) the maximum height would be 1700. So ... a reasonable range would be from 0 to 1700.